OpenStudy (anonymous):

Determine the focus and directrix of the given parabola, and sketch the graph: x^2 = -4y

OpenStudy (anonymous):

bring the equation to the standard form and get the directrix and focus using their formulae

OpenStudy (anonymous):

it is (0,0) and y = 0

OpenStudy (anonymous):

x^2 = -4y (1) First rearrange the equation into the standard form y=something y=(x^2) (2) ---- (-4) Now using (3) the standard derivation from the distance formula y=x^2 (3) --- 4(p) substitute the values from (2) so that 4xp = -4 y=x^2 --- 4(-1) so 4p = -4: with p=-1 p=-1 Directrix is -p so -(-1) is 1 so curve points downwards symetrical at 0 and focus is -1